# Interaction of the Primitive Equations with Sea Ice Dynamics

### v3

@@ 6. Primitive Equations on Cylindrical Domains
In this section, we collect properties of the primitive equations (4.1) on cylindrical domains $\Omega:=G\times(a,b)$, where $G=(0,1)\times(0,1)$ and $-\infty<a<b<\infty$. Apart from recalling that the hydrostatic Stokes operator admits a bounded $\mathcal{H}^{\infty}$-calculus, we provide estimates of the bilinear term appearing in the primitive equations tailored to our setting. Throughout this section, we denote by $v=(v_{1},v_{2})$ the principal variable corresponding to the horizontal velocity.
@@ 6.1. Bounded $\mathcal{H}^{\infty}$-calculus of the hydrostatic Stokes operator
[…]
Let $q\in(1,\infty)$ and $\mu\geq 0$. Then $-{\mathrm{A}}_{\mathrm{b.c.}}+\mu$ has a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)$ with $\Phi_{-{\mathrm{A}}_{\mathrm{b.c.}}+\mu}^{\infty}=0$ provided $\mu>0$. In the case $\Gamma_{\mathrm{D}}\neq\emptyset$, the above assertion is also valid for $\mu=0$.
Choosing $a=\kappa_{2}$, $b={h}_{\mathrm{atm}}$ and the boundary conditions (3.5), we obtain ${\mathrm{A}}_{\mathrm{b.c.}}=\mathrm{A}^{\mathrm{atm}}$, while the choice $a=-{h}_{\mathrm{ocn}}$, $b=0$ and the boundary conditions as in (4.3) lead to ${\mathrm{A}}_{\mathrm{b.c.}}=\mathrm{A}^{\mathrm{ocn}}_{0}$. Hence, we conclude the following from Subsection 6.1 and from results in [16, Section 3].
[…]
@@
    the operator $-\mathrm{A}^{\mathrm{atm}}+\mu$ admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})$ with $\Phi_{\mathrm{A}^{\mathrm{atm}}}^{\infty}=0$, and
@@
    the operator $-\mathrm{A}^{\mathrm{ocn}}_{0}$ admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})$ with $\Phi_{\mathrm{A}^{\mathrm{ocn}}_{0}}^{\infty}=0$, and it holds that $0\in\rho(-\mathrm{A}^{\mathrm{ocn}}_{0})$.
@@ 6.2. Estimates of the bilinearity
We recall the bilinearity $\mathrm{F}$ from (4.2). As a preparation, we briefly introduce anisotropic function spaces, see also [26, Section 5]. For $s,t\geq 0$ and $1\leq p,q\leq\infty$, we consider $\mathrm{W}_{z}^{{r},{q}}\mathrm{W}_{xy}^{{s},{p}}:=\mathrm{W}^{{r},q}((a,b);\mathrm{W}^{{s},p}(G))$ equipped with the norms $\|v\|_{\mathrm{W}_{z}^{{r},{q}}\mathrm{W}_{xy}^{{s},{p}}}=\left\|\|v(\cdot,z)\|_{\mathrm{W}^{{s},p}(G)}\right\|_{\mathrm{W}^{{r},q}(a,b)}$, so they become Banach spaces. The same remains valid when considering Bessel potential spaces instead of Sobolev spaces. First, employing Hölder’s inequality independently with respect to $z$ and $(x,y)$, we derive that for $p$, $p_{1}$, $p_{2}$ and $q$, $q_{1}$, $q_{2}$ such that $\nicefrac{{1}}{{p_{1}}}+\nicefrac{{1}}{{p_{2}}}=\nicefrac{{1}}{{p}}$ and $\nicefrac{{1}}{{q_{1}}}+\nicefrac{{1}}{{q_{2}}}=\nicefrac{{1}}{{q}}$, it holds that
@@
[…]
@@
The next result provides analogous estimates as in [26, Lemma 5.1] or [17, Lemma 6.1] for our setting.
[…]
@@
so it suffices to prove (a). As $\mathcal{P}\colon\mathrm{L}^{q}(\Omega)^{2}\to{\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)$ is bounded and $\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(\Omega)^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)\subset\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(\Omega)^{2}$ holds true, it is sufficient to bound the $\mathrm{L}^{q}(\Omega)$-norms of $({v}\cdot\nabla_{\mathrm{H}}){v}$ and $w({v})\cdot\partial_{z}{v}$ separately by $C\|v\|_{\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(\Omega)}^{2}$ for some $C>0$. Next, we deduce from $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$ the existence of $\eta>0$ such that $2-\nicefrac{{2}}{{p}}-\eta-\nicefrac{{3}}{{q}}\geq-\nicefrac{{3}}{{3q}}$ and $2-\nicefrac{{2}}{{p}}-\eta-\nicefrac{{3}}{{q}}\geq 1-\nicefrac{{2}}{{q}}$. It then follows by [49, Theorem 4.6.1] that
@@
where we used that $\mathrm{H}^{1,\nicefrac{{3q}}{{2}}}(\Omega)\hookrightarrow\mathrm{W}^{1,\nicefrac{{3q}}{{2}}}(\Omega)$ by virtue of $q\geq 2$, see e. g. [49, Chapter 2]. Combining the embeddings from (6.2) with Hölder’s inequality, we obtain
@@
The condition on $p$ and $q$ yields $1-\nicefrac{{2}}{{p}}-2\eta-\nicefrac{{2}}{{q}}\geq-\nicefrac{{2}}{{2q}}$ for some $\eta>0$, so [49, Theorem 4.6.1] implies
@@ 7. The Stationary Hydrostatic Stokes Problem
(7.1) $\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}&=\varphi,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{\Gamma_{b}},\end{aligned}\right.$
@@ 7.1. The hydrostatic Dirichlet operator
We start by solving (7.1) for smooth functions on the boundary. We need the following auxiliary result which guarantees the existence of an extension with vertical average 00. To this end, we construct a smooth solution to the associated extension problem
(7.2) $\overline{g}&=0,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}g&=\varphi,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}g&=0,&&\text{ on }{\Gamma_{b}}.\end{aligned}\right.$
@@
Note that ${\Omega}_{\mathrm{ocn}}=G\times(-{h}_{\mathrm{ocn}},0)$. The approach for $g$ is a splitting into the horizontal and the vertical part $g(x_{\mathrm{H}},z)=r(z)\cdot\varphi(x_{\mathrm{H}})$. Therefore, the problem (7.2) reduces to the construction of a smooth solution $r$ to the one-dimensional extension problem with mean value 00 given by
(7.3) $\overline{r}&=0,&&\text{ on }{(-{h}_{\mathrm{ocn}},0)},\\ r(0)&=1,&&\\ r(-{h}_{\mathrm{ocn}})&=0.&&\end{aligned}\right.$
@@
The regularity and average 00 of the solution $g$ of (7.2) ensure $g\in D(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})$, so $v_{g}:={v}_{\mathrm{ocn}}-g\in D(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})$ satisfies
(7.4) $\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}v_{g}&=f,&&\text{on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}v_{g}&=0,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}v_{g}&=0,&&\text{ on }{\Gamma_{b}},\end{aligned}\right.$
@@
From [3, Theorem VIII.1.3.1], establishing the result for the half-space, and the observation that horizontally periodic functions can be extended periodically onto a layer and then cut off, we deduce for $r>1+\nicefrac{{1}}{{q}}$ the embedding $\mathrm{H}_{\mathrm{per}}^{{r},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\hookrightarrow D(\partial_{\nu}^{q})$. The claim is then a consequence of Subsection 6.1, implying that $-\mathrm{A}^{\mathrm{ocn}}_{0}$ has bounded imaginary powers, see also the relation (8.7). ∎
@@
(7.6) $\Delta{v}_{\mathrm{ocn}}&=\nabla_{\mathrm{H}}\pi,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}&=\varphi,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{\Gamma_{b}}.\end{aligned}\right.$
@@
Assume now $s\geq 2-\nicefrac{{1}}{{q}}$ and $\varphi\in\mathrm{H}_{\mathrm{per}}^{{s},q}(G)^{2}$. By the above argument, we have ${v}_{\mathrm{ocn}}\in\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})$. Using the hydrostatic solenoidal condition, we obtain by a direct calculation similar to [16, (4.3)] that the pressure term is given by
@@
Assume that the a-priori bound still holds for $s-1$. Replacing the Calderón-Zygmund a-priori estimate (7.8) by its higher-order analogue and using the higher-order bounds for the normal derivative and the fact that the pseudo-differential operator $\nabla_{\mathrm{H}}\Delta_{\mathrm{H}}^{-1}\mathrm{div}_{\mathrm{H}}\,$ of order 00 is bounded on $\mathrm{H}_{\mathrm{per}}^{{s},q}(G)^{2}$, we deduce the induction step from the same argument as the induction base.
@@
$\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}{v}_{\mathrm{ocn}}&=f,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}&=\varphi,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{\Gamma_{b}},\end{aligned}\right.$
@@
[…]
Analogous results about stationary hydrostatic Stokes problems for pure Dirichlet or pure Neumann boundary conditions can be shown by the same technique. First of all, note that it is sufficient to consider only inhomogeneous boundary conditions, since the general case follows from linearity of the problem. The case of pure (inhomogeneous) Neumann boundary conditions is treated in [7, Section 3].

### v4

@@ 6. Primitive Equations on Cylindrical Domains
In this section, we collect properties of the primitive equations (4.1) on cylindrical domains $\Omega:=G\times(a,b)$, where $G=(0,1)\times(0,1)$ and $\infty<a<b<\infty$. Apart from recalling that the hydrostatic Stokes operator admits a bounded $\mathcal{H}^{\infty}$-calculus, we provide estimates of the bilinear term appearing in the primitive equations tailored to our setting. Throughout this section, we denote by $v=(v_{1},v_{2})$ the principal variable corresponding to the horizontal velocity.
@@ 6.1. Bounded $\mathcal{H}^{\infty}$-calculus of the hydrostatic Stokes operator
Proposition 6.1 (Theorem 3.1 in [17]).
Let $q\in(1,\infty)$ and $\mu\geq 0$. Then ${\mathrm{A}}_{\mathrm{b.c.}}+\mu$ has a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)$ with $\Phi_{-{\mathrm{A}}_{\mathrm{b.c.}}+\mu}^{\infty}=0$ provided $\mu>0$. In the case $\Gamma_{\mathrm{D}}\neq\emptyset$, the above assertion is also valid for $\mu=0$.
Choosing $a=\kappa_{2}$, $b={h}_{\mathrm{atm}}$ and the boundary conditions (3.5), we obtain ${\mathrm{A}}_{\mathrm{b.c.}}=\mathrm{A}^{\mathrm{atm}}$, while the choice $a=-{h}_{\mathrm{ocn}}$, $b=0$ and the boundary conditions as in (4.3) lead to ${\mathrm{A}}_{\mathrm{b.c.}}=\mathrm{A}^{\mathrm{ocn}}_{0}$. Hence, we conclude the following from Subsection 6.1 and from results in [17, Section 3].
Corollary 6.2.
@@
    the operator $\mathrm{A}^{\mathrm{atm}}+\mu$ admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{atm}})$ with $\Phi_{\mathrm{A}^{\mathrm{atm}}}^{\infty}=0$, and
@@
    the operator $\mathrm{A}^{\mathrm{ocn}}_{0}$ admits a bounded $\mathcal{H}^{\infty}$-calculus on ${\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})$ with $\Phi_{\mathrm{A}^{\mathrm{ocn}}_{0}}^{\infty}=0$, and it holds that $0\in\rho(-\mathrm{A}^{\mathrm{ocn}}_{0})$.
@@ 6.2. Estimates of the bilinearity
We recall the bilinearity $\mathrm{F}$ from (4.2). As a preparation, we briefly introduce anisotropic function spaces, see also [25, Section 5]. For $s,t\geq 0$ and $1\leq p,q\leq\infty$, we consider $\mathrm{W}_{z}^{{r},{q}}\mathrm{W}_{xy}^{{s},{p}}:=\mathrm{W}^{{r},q}((a,b);\mathrm{W}^{{s},p}(G))$ equipped with the norms $\|v\|_{\mathrm{W}_{z}^{{r},{q}}\mathrm{W}_{xy}^{{s},{p}}}=\left\|\|v(\cdot,z)\|_{\mathrm{W}^{{s},p}(G)}\right\|_{\mathrm{W}^{{r},q}(a,b)}$, so they become Banach spaces. The same remains valid when considering Bessel potential spaces instead of Sobolev spaces. First, employing Hölder’s inequality independently with respect to $z$ and $(x,y)$, we derive that for $p$, $p_{1}$, $p_{2}$ and $q$, $q_{1}$, $q_{2}$ such that $\nicefrac{{1}}{{p_{1}}}+\nicefrac{{1}}{{p_{2}}}=\nicefrac{{1}}{{p}}$ and $\nicefrac{{1}}{{q_{1}}}+\nicefrac{{1}}{{q_{2}}}=\nicefrac{{1}}{{q}}$, it holds that
@@
$\mathrm{W}_{z}^{{r},{q}}\mathrm{W}_{xy}^{{s},{p}}\hookrightarrow\mathrm{W}_{z}^{{r_{1}},{q_{1}}}\mathrm{W}_{xy}^{{s},{p}}$ $\text{ for }\mathrm{W}_{z}^{{r},{q}}(a,b)\hookrightarrow\mathrm{W}_{z}^{{r_{1}},{q_{1}}}(a,b),\text{ and }$
$\mathrm{W}_{z}^{{r},{q}}\mathrm{W}_{xy}^{{s},{p}}\hookrightarrow\mathrm{W}_{z}^{{r},{q}}\mathrm{W}_{xy}^{{s_{1}},{p_{1}}}$ $\text{ for }\mathrm{W}_{xy}^{{s},{p}}(G)\hookrightarrow\mathrm{W}_{xy}^{{s_{1}},{p_{1}}}(G).$
@@
The next result provides analogous estimates as in [25, Lemma 5.1] or [18, Lemma 6.1] for our setting.
Proposition 6.3.
@@
so it suffices to prove (a). As $\mathcal{P}\colon\mathrm{L}^{q}(\Omega)^{2}\to{\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)$ is bounded and $\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(\Omega)^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}(\Omega)\subset\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(\Omega)^{2}$ holds true, it is sufficient to bound the $\mathrm{L}^{q}(\Omega)$-norms of $({v}\cdot\nabla_{\mathrm{H}}){v}$ and $w({v})\cdot\partial_{z}{v}$ separately by $C\|v\|_{\mathrm{B}^{{2-\nicefrac{{2}}{{p}}}}_{{q}{p}}(\Omega)}^{2}$ for some $C>0$. Next, we deduce from $\nicefrac{{1}}{{p}}+\nicefrac{{1}}{{q}}<\nicefrac{{1}}{{2}}$ the existence of $\eta>0$ such that $2-\nicefrac{{2}}{{p}}-\eta-\nicefrac{{3}}{{q}}\geq-\nicefrac{{3}}{{3q}}$ and $2-\nicefrac{{2}}{{p}}-\eta-\nicefrac{{3}}{{q}}\geq 1-\nicefrac{{2}}{{q}}$. It then follows by [38, Theorem 4.6.1] that
@@
where we used that $\mathrm{H}^{1,\nicefrac{{3q}}{{2}}}(\Omega)\hookrightarrow\mathrm{W}^{1,\nicefrac{{3q}}{{2}}}(\Omega)$ by virtue of $q\geq 2$, see e. g. [38, Chapter 2]. Combining the embeddings from (6.2) with Hölder’s inequality, we obtain
@@
The condition on $p$ and $q$ yields $1-\nicefrac{{2}}{{p}}-2\eta-\nicefrac{{2}}{{q}}\geq-\nicefrac{{2}}{{2q}}$ for some $\eta>0$, so [38, Theorem 4.6.1] implies
@@ 7. The Stationary Hydrostatic Stokes Problem
(7.1) $\left\{\begin{aligned} \mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}&=\varphi,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{\Gamma_{b}},\end{aligned}\right$.
@@ 7.1. The hydrostatic Dirichlet operator
We start by solving (7.1) for smooth functions on the boundary. We need the following auxiliary result which guarantees the existence of an extension with vertical average 0. To this end, we construct a smooth solution to the associated extension problem
[…]
@@
Note that ${\Omega}_{\mathrm{ocn}}=G\times(-{h}_{\mathrm{ocn}},0)$. The approach for $g$ is a splitting into the horizontal and the vertical part $g(x_{\mathrm{H}},z)=r(z)\cdot\varphi(x_{\mathrm{H}})$. Therefore, the problem (7.2) reduces to the construction of a smooth solution $r$ to the one-dimensional extension problem with mean value 0 given by
[…]
@@
The regularity and average 0 of the solution $g$ of (7.2) ensure $g\in D(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})$, so $v_{g}:={v}_{\mathrm{ocn}}-g\in D(\mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}})$ satisfies
(7.4) $\left\{\begin{aligned} \mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}v_{g}&=f,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}v_{g}&=0,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}v_{g}&=0,&&\text{ on }{\Gamma_{b}},\end{aligned}\right$.
@@
From [3, Theorem VIII.1.3.1], establishing the result for the half-space, and the observation that horizontally periodic functions can be extended periodically onto a layer and then cut off, we deduce for $r>1+\nicefrac{{1}}{{q}}$ the embedding $\mathrm{H}_{\mathrm{per}}^{{r},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})\hookrightarrow D(\partial_{\nu}^{q})$. The claim is then a consequence of Subsection 6.1, implying that $\mathrm{A}^{\mathrm{ocn}}_{0}$ has bounded imaginary powers, see also the relation (8.7). ∎
@@
(7.6) $\left\{\begin{aligned} \Delta{v}_{\mathrm{ocn}}&=\nabla_{\mathrm{H}}\pi,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}&=\varphi,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{\Gamma_{b}}.\end{aligned}\right$.
@@
Assume now $s\geq 2-\nicefrac{{1}}{{q}}$ and $\varphi\in\mathrm{H}_{\mathrm{per}}^{{s},q}(G)^{2}$. By the above argument, we have ${v}_{\mathrm{ocn}}\in\mathrm{H}_{\mathrm{per}}^{{2},q}({\Omega}_{\mathrm{ocn}})^{2}\cap{\mathrm{L}}_{\overline{\sigma}}^{q}({\Omega}_{\mathrm{ocn}})$. Using the hydrostatic solenoidal condition, we obtain by a direct calculation similar to [17, (4.3)] that the pressure term is given by
@@
Assume that the a-priori bound still holds for s1s-1. Replacing the Calderón-Zygmund a-priori estimate (7.8) by its higher-order analogue and using the higher-order bounds for the normal derivative and the fact that the pseudo-differential operator $\nabla_{\mathrm{H}}\Delta_{\mathrm{H}}^{-1}\mathrm{div}_{\mathrm{H}}\,$ of order 0 is bounded on $\mathrm{H}_{\mathrm{per}}^{{s},q}(G)^{2}$, we deduce the induction step from the same argument as the induction base.
@@
$\left\{\begin{aligned} \mathrm{A}^{\mathrm{ocn}}_{\mathrm{m}}{v}_{\mathrm{ocn}}&=f,&&\text{ on }{{\Omega}_{\mathrm{ocn}}},\\ \mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}&=\varphi,&&\text{ on }{\Gamma_{o}},\\ \mathrm{tr}_{\Gamma_{b}}{v}_{\mathrm{ocn}}&=0,&&\text{ on }{\Gamma_{b}},\end{aligned}\right$.
@@
Remark 7.8.
Analogous results about stationary hydrostatic Stokes problems for pure Dirichlet or pure Neumann boundary conditions can be shown by the same technique. First of all, note that it is sufficient to consider only inhomogeneous boundary conditions, since the general case follows from linearity of the problem. The case of pure (inhomogeneous) Neumann boundary conditions is treated in [5, Section 3].